The following contingency table presents the analysis of 300 persons according to
hair colour and eye colour:
Hair Colour Eye Colour
Blue Grey Brown
Burgundy 30 10 40
Brown 40 20 40
Black 50 30 40

Test the hypothesis that there is an association between hair colour and eye colour at
5% level of significance
1
Expert's answer
2020-06-01T16:54:14-0400

H0 : Hair and eye colour are independent

H1 : Hair and eye colour are associated

χ2="\sum"ri=1 "\sum" cj=1=(Oij-Eij)2/Eij

Eij=RiCj/N, Ri - sum of i-th row, Cj - sum of j-th column, N - total sample size

E11=(80*120)/300=32, (O11-E11)2=4

E12=(80*60)/300=16, (O12-E12)2=36

E13=(80*120)/300=32, (O13-E13)2=64

E21=(100*120)/300=40, (O21-E21)2=0

E22=(100*60)/300=20, (O22-E22)2=0

E23=(100*120)/300=40, (O23-E23)2=0

E31=(120*120)/300=48, (O31-E31)2=4

E32=(120*60)/300=24, (O32-E32)2=36

E33=(120*120)/300=48, (O33-E33)2=64

χ2=7.29

degrees of freedom: (r-1)(c-1)=4

the critical value of chi square with 4df at 5% level significance: 9.488

χ2<critical value, so we accept H0 ,so hair colour is independent of eye colour